the length of a rectangle is four more than one-half the width of the rectangle. the perimeter of the rectangle is 50. The area of the rectangle is _______square units.
the length of the diagonal of the rectangle, to the nearest tenth is ________units.

Answer by stanbon(75887)

the length of a rectangle is four more than one-half the width of the rectangle.
------------------------------
Let width be "x".
Then the length is "(1/2)x+4"
------------------------------
the perimeter of the rectangle is 50.
P + 2(length + width)
50 = 2[(1/2)x+4 + x]
25 = (3/2)x + 4
(3/2)x = 21
x = 14 (width)
(1/2)x+4 = 7+4 = 11 (length)
---------------------------------
The area of the rectangle is 14*11 = 154 square units.
the length of the diagonal of the rectangle,
to the nearest tenth is sqrt(21^2+11^2) = sqrt(562) = 23.71 units.
===================================================================
Cheers,
Stan H.